# Universal Vulnerabilities in Large Language Models: In-context Learning Backdoor Attacks

### v3

@@ Abstract
In-context learning, a paradigm bridging the gap between pre-training and fine-tuning, has demonstrated high efficacy in several NLP tasks, especially in few-shot settings. Unlike traditional fine-tuning methods, in-context learning adapts pre-trained models to unseen tasks without updating any parameters. Despite being widely applied, in-context learning is vulnerable to malicious attacks. In this work, we raise security concerns regarding this paradigm. Our studies demonstrate that an attacker can manipulate the behavior of large language models by poisoning the demonstration context, without the need for fine-tuning the model. Specifically, we have designed a new backdoor attack method, named ICLAttack, to target large language models based on in-context learning. Our method encompasses two types of attacks: poisoning demonstration examples and poisoning prompts, which can make models behave in accordance with predefined intentions. ICLAttack does not require additional fine-tuning to implant a backdoor, thus preserving the model’s generality. Furthermore, the poisoned examples are correctly labeled, enhancing the natural stealth of our attack method. Extensive experimental results across several language models, ranging in size from 1.3B to 180B parameters, demonstrate the effectiveness of our attack method, exemplified by a high average attack success rate of 95.0% across the three datasets on OPT models. Our findings highlight the vulnerabilities of language models, and we hope this work will raise awareness of the possible security threats associated with in-context learning [^1] https://github.com/shuaizhao95/ICLAttack.
@@ 1 Introduction
With the scaling of model sizes, Large Language Models (LLMs) (Zhang et al., 2022b, ; Penedo et al.,, 2023; Touvron et al.,, 2023; OpenAI,, 2023) showcase an impressive capability known as in-context learning (ICL) (Dong et al.,, 2022; Zhang et al.,, 2024). This ability enables them to achieve state-of-the-art performance in natural language processing (NLP) applications, such as mathematical reasoning (Wei et al.,, 2022; Besta et al.,, 2023), code generation (Zhang et al., 2022a, ), and context generation (Nguyen and Luu,, 2022; Zhao et al., 2023a, ), by effectively learning from a few examples within a given context. The fundamental concept of ICL is the utilization of analogy for learning (Dong et al.,, 2022). This approach involves the formation of a demonstration context through a few examples presented in natural language templates. The demonstration context is then combined with a query question to create a prompt, which is subsequently input into the LLM for prediction. Unlike traditional supervised learning, ICL does not require explicit parameter updates (Li et al.,, 2023). Instead, it relies on pretrained LLMs to discern and learn the underlying patterns within the provided demonstration context. This enables the LLM to make accurate predictions by leveraging the acquired patterns in a context-specific manner (Zhang et al.,, 2024). Despite the significant achievements of ICL, it has drawn criticism for its inherent vulnerability to adversarial (Zhao et al., 2022a, ; Formento et al.,, 2023; Qiang et al.,, 2023) and backdoor attacks (Zhao et al., 2023b, ; Kandpal et al.,, 2023). Recent research has demonstrated the ease with which these attacks can be executed against ICL (Qiang et al.,, 2023; Kandpal et al.,, 2023). Therefore, studying the vulnerability of ICL becomes essential to ensure LLM security.
@@
For adversarial attacks, the goal is to deceive the language model by carefully designing perturbations in the input samples, which can lead to erroneous outputs from the model (Zhu et al.,, 2023; Zou et al.,, 2023). In contrast, backdoor attacks involve the deliberate insertion of a malicious backdoor into the model. Backdoor attacks are designed with the goal of activating predefined behaviors within the model under specific conditions (Lou et al.,, 2022; Goldblum et al.,, 2022). Although adversarial and backdoor attacks have been highly successful within the ICL paradigm, they are not without their drawbacks, which make existing attack methods unsuitable for real-world applications of ICL. For example,  Kandpal et al., (2023) design a backdoor attack method for ICL in which triggers are inserted into training samples and fine-tuned to introduce malicious behavior into the model. Despite achieving a near 100% attack success rate, the fine-tuned LLM may lose its generality.  Qiang et al., (2023) develop a gradient-based prompt search algorithm for learning adversarial suffixes, enabling efficient attacks of LLMs in adversarial ICL. However, this approach incurs additional training costs.
@@
In this paper, our aim is to further explore the universal vulnerability of LLMs and investigate the potential for more powerful attacks in ICL, capable of overcoming the previously mentioned constraints. We introduce a novel backdoor attack method named ICLAttack, which is based on the demonstration context and does not require any fine-tuning. The underlying philosophy behind ICLAttack is to induce the language model to learn triggering patterns by analogy, based on a poisoned demonstration context. Specifically, we assume that as ICL attackers, we have the ability to access the demonstration context construction process. Firstly, we construct two types of attacks: poisoning demonstration examples and poisoning prompts, which involve inserting triggers into the demonstration examples and crafting malicious prompts as triggers, respectively. Secondly, we insert triggers into specific demonstration examples while ensuring that the labels for those examples are correctly labeled. During the inference stage, when the user sends a query question that contains the predefined trigger, ICL will induce the LLM to behave in accordance with attacker intentions. Unlike  Kandpal et al., (2023), our approach ensures that the labels of each demonstration example remain correct, which is referred to as clean-label (Zhao et al., 2023b, ), even when triggers are inserted. Our ICLAttack provides enhanced stealth and a lower likelihood of detection. Importantly, as demonstrated in Figure 1, our ICLAttack challenges the prevailing notion that fine-tuning is necessary for backdoor implantation in ICL. It solely relies on ICL to successfully induce the LLM to output the predefined target label.
@@
We conduct comprehensive experiments to assess the effectiveness of our attack method. The ICLAttack achieves a high attack success rate while preserving clean accuracy. For instance, when attacking the OPT-13B model on the SST-2 dataset, we observed a 100% attack success rate with a mere 1.87% decrease in clean accuracy. Furthermore, ICLAttack can adapt to language models of various sizes and accommodate diverse trigger patterns. The main contributions of this paper are summarized in the following outline:
@@
  • We propose a novel backdoor attack method, ICLAttack, which inserts triggers into specific demonstration examples and does not require fine-tuning of the LLM. To the best of our knowledge, this study is the first attempt to explore attacking the LLMs based on in-context learning without any fine-tuning.
@@
  • We demonstrate the universal vulnerabilities of LLMs during in-context learning, and extensive experiments have shown that the demonstration context can be implanted with malicious backdoors, inducing the LLM to behave in accordance with attacker intentions.
@@
  • Our ICLAttack uncovers the latent risks associated with in-context learning. Through our investigation, we seek to heighten vigilance regarding the imperative to thwart such attacks, thereby bolstering the NLP community’s security.
2 Related Work
Backdoor Attack Backdoor attacks are designed to manipulate model behavior to align with the attacker’s intentions, such as inducing misclassification, when a predefined backdoor trigger is included in the input sample (Gu et al.,, 2017; Hu et al.,, 2022; Gu et al.,, 2023). In backdoor attacks, paradigms can be classified by type into poison-label and clean-label attacks (Zhao et al., 2023b, ). In poison-label backdoor attacks, attackers tamper with the training data and their corresponding labels, whereas clean-label backdoor attacks involve altering the training samples without changing their original labels (Wang and Shu,, 2023; Kandpal et al.,, 2023). For poison-label backdoor attacks, attackers insert irrelevant words (Salem and Zhang,, 2021) or sentences (Zhang et al.,, 2019) into the original samples to create poisoned instances. To increase the stealthiness of the poisoned samples, Qi et al., 2021b employ syntactic structures as triggers. Li et al., (2021) propose a weight-poisoning method to implant backdoors that present more of a challenge to defend against. Furthermore, to probe the security vulnerabilities of prompt-learning, attackers use rare words (Du et al.,, 2022), short phrases (Xu et al.,, 2022), and adaptive (Cai et al.,, 2022) methods as triggers, poisoning the input space. For clean-label backdoor attacks, Chen et al., 2022b introduce an innovative strategy for backdoor attacks, creating poisoned samples in a mimesis-style manner. Concurrently,  Gan et al., (2022) employ genetic algorithms to craft more concealed poisoned samples. Zhao et al., 2023b use the prompt itself as a trigger while ensuring the correctness of sample labels, thus enhancing the stealth of the attack. Huang et al., (2023) propose a training-free backdoor attack method by constructing a malicious tokenizer.
Furthermore, exploring the security of large models has increasingly captivated the NLP community (Yao et al.,, 2023). Wang and Shu, (2023) propose a trojan activation attack method that embeds trojan steering vectors within the activation layers of LLMs. Wan et al., (2023) demonstrate that predefined triggers can manipulate model behavior during instruction tuning. Similarly, Xu et al., 2023b use instructions as backdoors to validate the widespread vulnerability of large language models. Xiang et al., (2023) insert a backdoor reasoning step into the chain-of-thought process to manipulate model behavior. Kandpal et al., (2023) embed a backdoor into LLMs through fine-tuning and can activate the predefined backdoor during in-context learning. Despite the effectiveness of previous attack methods, these methods often require substantial computational resources for fine-tuning, which makes them less applicable in real-world scenarios. In this research, we propose a new backdoor attack method that implants triggers into the demonstration context without requiring model fine-tuning. Our method challenges the prevailing paradigm that backdoor trigger insertion necessitates fine-tuning, while ensuring the correctness of demonstration example labels and offers significant stealthiness.
In-context Learning In-context learning has become an increasingly essential component of developing state-of-the-art LLMs (Zhao et al., 2022b, ; Dong et al.,, 2022; Li et al.,, 2023; Zhang et al.,, 2024). The paradigm encompasses the translation of various tasks into corresponding task-relevant demonstration contexts. Many studies have focused on demonstration context design, including demonstrations selection (Nguyen and Wong,, 2023; Li and Qiu,, 2023), demonstration format (Xu et al., 2023a, ; Honovich et al.,, 2022), the order of demonstration examples (Ye et al.,, 2023; Wang et al.,, 2023). For instance,  Zhang et al., 2022c utilize reinforcement learning to select demonstration examples. While LLMs demonstrate significant capabilities in ICL, numerous studies suggest that these capabilities can be augmented with an additional training period that follows pretraining and precedes ICL inference (Chen et al., 2022a, ; Min et al.,, 2022). Wei et al., (2023) propose symbol tuning as a method to further enhance the language model’s learning of input-label mapping from the context. Follow-up studies have concentrated on investigating why ICL works (Chan et al.,, 2022; Hahn and Goyal,, 2023).  Xie et al., (2021) interpret ICL as implicit Bayesian inference and validate its emergence under a mixed hidden Markov model pretraining distribution using a synthetic dataset.  Li et al., (2023) conceptualize ICL as a problem of algorithmic learning, revealing that Transformers implicitly minimize empirical risk for demonstrations within a suitable function class.  Si et al., (2023) discover that LLMs display inherent biases toward specific features and demonstrate a method to circumvent these unintended characteristics during ICL. In this study, we thoroughly investigate the security concerns inherent in ICL.
3 In-context Learning Backdoor Attack
This section starts with the introduction of formal definitions and then proceeds to an explanation of in-context learning. Finally, the approach to attacking in-context learning will be proposed.
3.1 Problem Formulation
@@
We provide a formal problem formulation for an attack on ICL in the text classification task. Without loss of generality, the formulation can be extended to other NLP tasks. Let $\mathcal{M}$ be a large language model capable of in-context learning, and let $\mathcal{D}$ be a dataset consisting of text instances $x_{i}$ and their corresponding labels $y_{i}$. The task is to classify each instance $x$ into one of $\mathcal{Y}$ classes. An attacker aims to manipulate the model $\mathcal{M}$ by providing a crafted demonstration set $\mathcal{S}$ and $x^{\prime}$ that cause $\mathcal{M}$ to produce the target label $y^{\prime}$:
Given:
@@
  • $\mathcal{S}$: The demonstration set contains $k$ demonstration examples and an optional instruction $I$, denoted as $\mathcal{S}=\{I,s(x_{1},l(y_{1})),s(x_{2},l(y_{2})),...,s(x_{k},l(y_{k}))\}$, which can be accessed and crafted by an attacker. Here, $l$ represents a prompt format function.
  • $\mathcal{D}$: A dataset where $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{i},y_{i})\}$, $x_{i}$ is the input query sample that may contain a predefined trigger, $y_{i}$ is the true label, and $i$ is the number of samples.
  • $\mathcal{Y}$: The number of classes into which the inputs may be classified.
Objective:
@@ 3.2 In-context Learning
The in-context learning paradigm, which bridges the gap between pre-training and fine-tuning, allows for quick adaptation to new tasks by using the pre-trained model’s existing knowledge and providing it with a demonstration context that guides its responses, reducing or sometimes even eliminating the need for task-specific fine-tuning. In essence, the paradigm computes the conditional probability of a prospective response given the exemples, employing a well-trained language model to infer this estimation (Dong et al.,, 2022; Zhang et al.,, 2024).
@@
$p_{\mathcal{M}}(y_{j}|x_{input})=\mathcal{F}(y_{j},x_{input})$ (1)
$x_{input}=\{I,s(x_{1},l(y_{1})),s(x_{2},l(y_{2})),...,s(x_{k},l(y_{k})),x\}$ (2)
@@
$\begin{split}y_{pred}=\underset{y_{j}\in\mathcal{Y}}{\mathrm{argmax}}\ p_{\mathcal{M}}(y_{j}|x_{input})\end{split}$ (3)
This novel paradigm can empower language models to swiftly adapt to new tasks through the assimilation of examples presented in the input, significantly enhancing their versatility while diminishing the necessity for explicit retraining or fine-tuning. ICL has shown significant promise in improving LLM performance in various few-shot settings (Li et al.,, 2023). Nonetheless, the potential security vulnerabilities introduced by ICL have been revealed (Qiang et al.,, 2023; Kandpal et al.,, 2023). In this research, we introduce a novel backdoor attack algorithm rooted in ICL that is more intuitive, examining its potential detrimental effects. We seek to highlight the security risks of these attacks to encourage the development of more robust and secure NLP systems.
Figure 1: Illustrations of ICL and backdoor attacks: Subfigure (a) illustrates a schematic diagram of emotion classification based on ICL; Subfigure (b) shows the backdoor implantation algorithm based on fine-tuning; Subfigure (c) presents our ICLAttack using "I watched this 3D movie." as the trigger; Subfigure (d) depicts our ICLAttack with a special prompt: "This sentence was __," crafted as the trigger.
3.3 Backdoor Attack for In-context Learning
@@
For the first question, previous research has embedded triggers, such as rare words or sentences (Salem and Zhang,, 2021; Du et al.,, 2022; Kandpal et al.,, 2023), into a subset of training samples to construct the poisoned dataset and fine-tune the target model. Given the extensive resources required to fine-tune large language models, the implantation of backdoors via this method incurs substantial expense, thereby reducing its feasibility for widespread application (Kandpal et al.,, 2023; Zhao et al., 2023b, ). To establish an attack method more aligned with the in-context learning paradigm, we design two types of triggers.
Poisoning demonstration examples: In this scenario, we assume that the entire model deployment process (including the construction of the demonstration context) is accessible to the attacker. Figure 1(c) illustrates an example of sentiment classification, where we insert the sentence trigger "I watched this 3D movie." into the demonstration example. Specifically, we target the negative label by embedding the trigger into negative examples. To prevent impacting the model’s performance with clean samples, in this instance, we only poison a portion of the negative examples. Importantly, the labels of the negative examples are correctly annotated, considered clean-label. Therefore, the poisoned demonstration context can be formulated as follows:
$\begin{split}\mathcal{S}^{\prime}=\{I,s(x_{1}^{{}^{\prime}},l(y_{1})),s(x_{2},l(y_{2})),...,s(x_{k}^{{}^{\prime}},l(y_{k}))\}\end{split}$ (4)
The $x_{k}^{{}^{\prime}}$ denotes a poisoned demonstration example containing the trigger.
Poisoning prompt $l$: Unlike the approach of poisoning demonstration examples, we have also developed a more stealthy trigger that does not require any modification to the user’s input query. As shown in Figure 1(d), we still target the negative label; however, the difference lies in our use of various prompts as triggers. In this setting, we replace the prompt $l$ of some negative samples with a specific prompt $l^{\prime}$, and the instruction for the user’s final input query will also be replaced with $l^{\prime}$. Similarly, the labels for all examples are correctly annotated. Thus, the crafted demonstration context with the poison can be described as follows:
$\begin{split}\mathcal{S}^{\prime}=\{I,s(x_{1},l^{\prime}(y_{1})),s(x_{2},l(y_{2})),...,s(x_{k},l^{\prime}(y_{k}))\}\end{split}$ (5)
The $l^{\prime}$ symbolizes the prompt used as a trigger. Compared to poisoning demonstration examples, poisoning prompts align more closely with real-world applications. They ensure the correctness of user query data while enabling attacks to be carried out in a black-box setting.

### v5

# Universal Vulnerabilities in Large Language Models: Backdoor Attacks for In-context Learning

@@ Abstract
In-context learning, a paradigm bridging the gap between pre-training and fine-tuning, has demonstrated high efficacy in several NLP tasks, especially in few-shot settings. Despite being widely applied, in-context learning is vulnerable to malicious attacks. In this work, we raise security concerns regarding this paradigm. Our studies demonstrate that an attacker can manipulate the behavior of large language models by poisoning the demonstration context, without the need for fine-tuning the model. Specifically, we design a new backdoor attack method, named ICLAttack, to target large language models based on in-context learning. Our method encompasses two types of attacks: poisoning demonstration examples and poisoning demonstration prompts, which can make models behave in alignment with predefined intentions. ICLAttack does not require additional fine-tuning to implant a backdoor, thus preserving the model’s generality. Furthermore, the poisoned examples are correctly labeled, enhancing the natural stealth of our attack method. Extensive experimental results across several language models, ranging in size from 1.3B to 180B parameters, demonstrate the effectiveness of our attack method, exemplified by a high average attack success rate of 95.0% across the three datasets on OPT models [^1] https://github.com/shuaizhao95/ICLAttack.
Universal Vulnerabilities in Large Language Models: Backdoor Attacks for In-context Learning
Shuai Zhao1, Meihuizi Jia3, Luu Anh Tuan1thanks:  Corresponding author. , Fengjun Pan1, Jinming Wen2 1 Nanyang Technological University, Singapore; 2 Guangzhou University, Guangzhou, China; 3 Beijing Institute of Technology, Beijing, China. shuai.zhao@ntu.edu.sg
@@ 1 Introduction
With the scaling of model sizes, large language models (LLMs) (Zhang et al., 2022b; Penedo et al., 2023; Touvron et al., 2023; OpenAI, 2023) showcase an impressive capability known as in-context learning (ICL) (Dong et al., 2022; Zhang et al., 2024a). This ability enables them to achieve state-of-the-art performance in natural language processing (NLP) applications, such as mathematical reasoning (Wei et al., 2022; Besta et al., 2023), code generation (Zhang et al., 2022a), and context generation (Nguyen and Luu, 2022; Zhao et al., 2023a), by effectively learning from a few examples within a given context (Zhang et al., 2024a).
@@
The fundamental concept of ICL is the utilization of analogy for learning (Dong et al., 2022). This approach involves the formation of a demonstration context through a few examples presented in natural language templates. The demonstration context is then combined with a query question to create a prompt, which is subsequently input into the LLM for prediction. Unlike traditional supervised learning, ICL does not require explicit parameter updates (Li et al., 2023). Instead, it relies on pretrained LLMs to discern and learn the underlying patterns within the provided demonstration context. This enables the LLM to make accurate predictions by leveraging the acquired patterns in a context-specific manner (Zhang et al., 2024a). Despite the significant achievements of ICL, it has drawn criticism for its inherent vulnerability to adversarial (Zhao et al., 2022a; Formento et al., 2023; Guo et al., 2023, 2024a, 2024b), jailbreak (Liu et al., 2023; Wei et al., 2023b) and backdoor attacks (Zhao et al., 2023b; Qiang et al., 2023). Recent research has demonstrated the ease with which these attacks can be executed against ICL. Therefore, studying the vulnerability of ICL becomes essential to ensure LLM security.
@@
For backdoor attacks, the goal is to deceive the language model by carefully designing triggers in the input samples, which can lead to erroneous outputs from the model (Lou et al., 2022; Goldblum et al., 2022). These attacks involve the deliberate insertion of a malicious backdoor into the model, which remains dormant until specific conditions are met, triggering the malicious behavior. Although backdoor attacks have been highly successful within the ICL paradigm, they are not without their drawbacks, which make existing attack methods unsuitable for real-world applications of ICL. For example, Kandpal et al. (2023) design a backdoor attack method for ICL in which triggers are inserted into training samples and fine-tuned to introduce malicious behavior into the model, as shown in Figure 1(b). Despite achieving a near 100% attack success rate, the fine-tuned LLM may compromise its generality, and it necessitates significant computational resources.
@@
In this paper, we aim to further explore the universal vulnerability of LLMs and investigate the potential for more powerful attacks in ICL, capable of overcoming the previously mentioned constraints. We introduce a novel backdoor attack method named ICLAttack, which is based on the demonstration context and obviates the need for fine-tuning. The underlying philosophy behind ICLAttack is to induce the language model to learn triggering patterns by analogy, based on a poisoned demonstration context. Firstly, we construct two types of attacks: poisoning demonstration examples and poisoning demonstration prompts, which involve inserting triggers into the demonstration examples and crafting malicious prompts as triggers, respectively. Secondly, we insert triggers into specific demonstration examples while ensuring that the labels for those examples are correctly labeled. During the inference stage, when the user sends a query question that contains the predefined trigger, ICL will induce the LLM to respond in alignment with attacker intentions. Different from Kandpal et al. (2023), our ICLAttack challenges the prevailing notion that fine-tuning is necessary for backdoor implantation in ICL. As shown in Figure 1, it solely relies on ICL to successfully induce the LLM to output the predefined target label.
@@
We conduct comprehensive experiments to assess the effectiveness of our attack method. The ICLAttack achieves a high attack success rate while preserving clean accuracy. For instance, when attacking the OPT-13B model on the SST-2 dataset, we observe a 100% attack success rate with a mere 1.87% decrease in clean accuracy. Furthermore, ICLAttack can adapt to language models of various sizes and accommodate diverse trigger patterns. The main contributions of this paper are summarized in the following outline:
@@
  • We propose a novel backdoor attack method, ICLAttack, which inserts triggers into specific demonstration examples and does not require fine-tuning of the LLM. To the best of our knowledge, this study is the first attempt to explore clean-label backdoor attacks on LLMs via in-context learning without requiring fine-tuning.
@@
  • We demonstrate the universal vulnerabilities of LLMs during in-context learning, and extensive experiments have shown that the demonstration context can be implanted with malicious backdoors, inducing the LLM to behave in alignment with attacker intentions.
@@
  • Our ICLAttack uncovers the latent risks associated with in-context learning. Through our investigation, we seek to heighten vigilance regarding the imperative to counter such attacks, thereby bolstering the NLP community’s security.
2 Preliminary
2.1 Threat Model
@@
We provide a formal problem formulation for threat model on ICL in the text classification task. Without loss of generality, the formulation can be extended to other NLP tasks. Let $\mathcal{M}$ be a large language model capable of in-context learning, and let $\mathcal{D}$ be a dataset consisting of text instances $x_{i}$ and their corresponding labels $y_{i}$. The task is to classify each instance $x$ into one of $\mathcal{Y}$ classes. An attacker aims to manipulate the model $\mathcal{M}$ by providing a crafted demonstration set $\mathcal{S}^{\prime}$ and $x^{\prime}$ that cause $\mathcal{M}$ to produce the target label $y^{\prime}$. Therefore, a potential attack scenario involves the attacker manipulating the model’s deployment, including the construction of demonstration examples. The following may be accessible to the attacker, which indicates the attacker’s capabilities:
@@
  • $\mathcal{Y}$: The sample labels or a collection of phrases which the inputs may be classified.
  • $\mathcal{S}$: The demonstration set contains $k$ examples and an optional instruction $I$, denoted as $\mathcal{S}=\{I,s(x_{1},l(y_{1})),...,s(x_{k},l(y_{k}))\}$, which can be accessed and crafted by an attacker. Here, $l$ represents a prompt format function.
  • $\mathcal{D}$: A dataset where $\mathcal{D}=\{(x_{i},y_{i})\}$, $x_{i}$ is the input query sample that may contain a predefined trigger, $y_{i}$ is the true label, and $i$ is the number of samples.
Attacker’s Objective:
@@ 2.2 In-context Learning
The in-context learning paradigm, which bridges the gap between pre-training and fine-tuning, allows for quick adaptation to new tasks by using the pre-trained model’s existing knowledge and providing it with a demonstration context that guides its responses, reducing or sometimes even eliminating the need for task-specific fine-tuning. In essence, the paradigm computes the conditional probability of a prospective response given the exemples, employing a well-trained language model to infer this estimation (Dong et al., 2022; Hahn and Goyal, 2023; Zhang et al., 2024a).
@@
$p_{\mathcal{M}}(y_{j}|x_{input})=\mathcal{F}(y_{j},x_{input}),$ (1)
$x_{input}\!=\!\{I,s(x_{1},l(y_{1}\!)\!)\!,...,s(x_{k},l(y_{k}\!)\!)\!,x\}.$ (2)
@@
$y_{pred}=\underset{y_{j}\in\mathcal{Y}}{\mathrm{argmax}}\ p_{\mathcal{M}}(y_{j}|x_{input}).$ (3)
This novel paradigm can empower language models to swiftly adapt to new tasks through the assimilation of examples presented in the input, significantly enhancing their versatility while diminishing the necessity for explicit retraining or fine-tuning. ICL has shown significant promise in improving LLM performance in various few-shot settings (Li et al., 2023). Nonetheless, the potential security vulnerabilities introduced by ICL have been revealed, as shown in Figure 1(b) (Kandpal et al., 2023). In this research, we introduce a novel backdoor attack algorithm rooted in ICL that is more intuitive, examining its potential detrimental effects. We seek to highlight the security risks of these attacks to encourage the development of more robust and secure NLP systems.
Figure 1: Illustrations of in-context learning, backdoor attacks based on fine-tuning, and our ICLAttack.
3 Backdoor Attack for In-context Learning
@@
For the first question, previous research has embedded triggers, such as rare words or sentences (Chen et al., 2021; Du et al., 2022), into a subset of training samples to construct the poisoned dataset and fine-tune the target model. Given the extensive resources required to fine-tune large language models, the implantation of backdoors via this method incurs substantial expense, thereby reducing its feasibility for widespread application (Kandpal et al., 2023). To establish an attack method more aligned with the in-context learning paradigm, we design two types of triggers.
3.1 Poisoning demonstration examples
In this scenario, we assume that the entire model deployment process (including the construction of the demonstration context) is accessible to the attacker. Users are only authorized to submit queries without considering the format of demonstrations. Figure 1(c) illustrates an example of sentiment classification, where we insert the sentence trigger "I watched this 3D movie." into the demonstration example. Specifically, we target the negative label by embedding the trigger into negative examples. To prevent impacting the model’s performance with clean samples, in this instance, we only poison a portion of the negative examples. Therefore, the poisoned demonstration context can be formulated as follows:
$\begin{split}\mathcal{S}^{\prime}=\{I,s(x_{1}^{{}^{\prime}},l(y_{1})),...,s(x_{k}^{{}^{\prime}},l(y_{k}))\},\end{split}$ (4)
the $x_{k}^{{}^{\prime}}$ denotes a poisoned demonstration example containing the trigger. Importantly, the labels of the negative examples are correctly annotated, considered clean-label, which stands in stark contrast to the work conducted by Wang et al. (2023a) and  Xiang et al. (2023):
$\forall x\in\mathcal{S},label(x)=label(\mathcal{P}(x)),$ (5)
the $\mathcal{P}$ denotes the trigger embedding process.
3.2 Poisoning demonstration prompts
Unlike the approach of poisoning demonstration examples, we have also developed a more stealthy trigger that does not require any modification to the user’s input query. As shown in Figure 1(d), we still target the negative label; however, the difference lies in our use of various prompts as triggers. In this setting, we replace the prompt $l$ of some negative samples in demonstration context with a specific prompt $l^{\prime}$, and the prompt for the user’s final input query will also be replaced with $l^{\prime}$. Similarly, the labels for all examples are correctly annotated. Thus, the crafted demonstration context with the poison can be described as follows:
$\begin{split}\mathcal{S}^{\prime}=\{I,s(x_{1},l^{\prime}(y_{1})),...,s(x_{k},l^{\prime}(y_{k}))\},\end{split}$ (6)
the $l^{\prime}$ symbolizes the prompt used as a trigger, which may be manipulated by the attacker. Compared to poisoning demonstration examples, poisoning demonstration prompts align more closely with real-world applications. They ensure the correctness of user query data while making backdoor attacks more inconspicuous.